OpenStudy (anonymous):

Write the ratio as a ratio of whole numbers in lowest terms $0.80 to $1.00

OpenStudy (anonymous):

\[\frac{ 0.80 }{ 1.00 }=\frac{ 80 }{100 }=\frac{ 4 }{5 }\]

OpenStudy (anonymous):

Write each rate as a fraction in lowest terms 40 feet in 30 seconds